Mean Free Path — Question 6

Section 19-5 · Physics pages 3122–3123

Question 6 Different stationary targets contribute additive collision hazards A fast probe travels through two types of effectively stationary scatterers. Their number densities are νA = 2.00 × 1020 and νB = 1.00 × 1020 m−3; cross sections are σA = 1.00 × 10−19 and σB = 3.00 × 10−19 m2. Independent hazards add: 1/λ = νAσA + νBσB. No identical-moving-gas √ 2 factor applies. stationary targets A + B Representative particles; dots do not depict the actual count or scale. (a) Find the mean free path and the probability that a collision, when it occurs, is with B. (b) Find survival through 3.00 cm. If B is removed, calculate the new mean free path and survival.
Show solutionHide solution
Question 6 - Solution Different stationary targets contribute additive collision hazards stationary targets A + B (a) Method and calculation hA = 20.0 m−1, hB = 30.0 m−1, λ = 1/(20 + 30) = .0200 m. P(B | collision) = hB/(hA + hB) = .600. (b) Interpretation and further calculation S(.0300) = e−50(.0300) = e−1.5 = .223130. λA only = 1/20 = .0500 m, SA = e−.600 = .548812. A less numerous species can dominate collisions when its cross section is larger. Physical check Hazards, rather than mean free paths, add for independent competing collision mechanisms.

Question 6

Different stationary targets contribute additive collision hazards

A fast probe travels through two types of effectively stationary scatterers. Their number densities are νA=2.00×1020\nu_A=2.00\times10^{20} and νB=1.00×1020m−3\nu_B=1.00\times10^{20}\,\mathrm{m^{-3}}; cross sections are σA=1.00×10−19\sigma_A=1.00\times10^{-19} and σB=3.00×10−19m2\sigma_B=3.00\times10^{-19}\,\mathrm{m^2}. Independent hazards add: 1/λ=νAσA+νBσB1/\lambda=\nu_A\sigma_A+\nu_B\sigma_B. No identical-moving-gas 2\sqrt2 factor applies.

Diagram shown in the original SVG worksheet.

  1. Find the mean free path and the probability that a collision, when it occurs, is with B.

  2. Find survival through 3.00 cm. If B is removed, calculate the new mean free path and survival.

Question 6 - Solution

Different stationary targets contribute additive collision hazards

Diagram shown in the original SVG worksheet.

(a) Method and calculation

hA=20.0m−1,hB=30.0m−1,λ=1/(20+30)=.0200m.h_A=20.0\,\mathrm{m^{-1}},\quad h_B=30.0\,\mathrm{m^{-1}},\quad\lambda=1/(20+30)=.0200\,\mathrm m.P(B∣collision)=hB/(hA+hB)=.600.P(B\mid\text{collision})=h_B/(h_A+h_B)=.600.

(b) Interpretation and further calculation

S(.0300)=e−50(.0300)=e−1.5=.223130.S(.0300)=e^{-50(.0300)}=e^{-1.5}=.223130.λAonly=1/20=.0500m,SA=e−.600=.548812.\lambda_{A\mathrm{\ only}}=1/20=.0500\,\mathrm m,\quad S_A=e^{-.600}=.548812.A less numerous species can dominate collisions when its cross section is larger.

Physical check

Hazards, rather than mean free paths, add for independent competing collision mechanisms.